Hi. I'm having trouble finding the surface area for my geometry homework. I need an answer in the next 20 mins.! PLEASE HELP!

One shape is a cube. The base is 7 cm.
One shape is a rectangular prisim. The base is 5 in. The face is 3 in. The height is 1 in.
One shape is a rectangular prisim. Base is 5 in. Face is 4 in. Height is 3 in.
One shape is a rectangular pyramid. Base 4.5 meters. Height 4.5 m. Face 6.5 m.
One shape is rectangular prisim. Base 4.5 cm. Face 3 cm. Height 3 cm.
One shape is rectangular pyramid. Base 2 in. Face 2 in. Height 5 in.
PLEASE DO THOSE PROBLEMS FOR ME! I THINK I GOT THEM ALL WRONG! THIS IS GRADED! I NEED ANSWERS NOW! DO IT FOR ME PLEASE! GIVE ME THE ANSWERS!!!!! (:

Answer by drk(1908)

I assume by base, face, height, you mean length, width, height.
----
One shape is a cube. The base is 7 cm.
answer
SA = 6s^2
SA = 6*7^2 = 6*49 = 494 cm^2
----
One shape is a rectangular prisim. The base is 5 in. The face is 3 in. The height is 1 in.
SA = 2*(lw + lh+ wh)
SA = 2*(5*3 + 5*1 +3*1)
SA = 2(15 + 5 + 3)
SA = 2*(23) = 46 in ^2
----
One shape is a rectangular prisim. Base is 5 in. Face is 4 in. Height is 3 in.
SA = 2*(lw + lh+ wh)
SA = 2*(5*4 + 5*3 +4*3)
SA = 2(20 + 15 + 12)
SA = 2*(47) = 94 in ^2
---
One shape is a rectangular pyramid. Base 4.5 meters. Height 4.5 m. Face 6.5 m.
SA = 2*(1/2)*sqrt(4.5^2-2.25^2) + 2*(1/2)*sqrt(4.5^2-3.25^2) + 4.5*6.5
SA = 36.2596 m^2
----
One shape is rectangular prisim. Base 4.5 cm. Face 3 cm. Height 3 cm.
SA = 2*(lw + lh+ wh)
SA = 2*(4.5*3 + 4.5*3 +3*3)
SA = 2(13.5 + 13.5 + 9)
SA = 2*(36) = 72 cm ^2
----
One shape is rectangular pyramid. Base 2 in. Face 2 in. Height 5 in.
SA = 2*(1/2)*sqrt(5^2-1^2) + 2*(1/2)*sqrt(5^2-1^2) + 2*5
SA = 19.7979 in^2